OpenStudy (anonymous):

Verify the identity: cos(x+pi/2) = -sin x I think the point of this question is to use the sum/difference formula.

jimthompson5910 (jim_thompson5910):

you would use the identity cos(A+B) = cos(A)cos(B) - sin(A)sin(B)

OpenStudy (anonymous):

I have this so far: cos x cos pi/2 - sin x sin pi/2 = -sin x

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

cos(pi/2) = ???

OpenStudy (anonymous):

I have no idea. I think it's 0 but I'm not entirely sure.

jimthompson5910 (jim_thompson5910):

yes, using the unit circle (get to know it or have it on a sheet of paper handy) you'll find that cos(pi/2) = 0 sin(pi/2) = 1

OpenStudy (anonymous):

Sooo, cos x (0) - sin x (1) = -sin x ?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

the cos(x) and 0 multiply to 0

jimthompson5910 (jim_thompson5910):

so the left half of cos x (0) - sin x (1) goes away

jimthompson5910 (jim_thompson5910):

leaving you with -sin(x) * 1 = -sin(x)

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

yw